OpenStudy (anonymous):

write y-10=2(x-8) in standard from

OpenStudy (hoblos):

-2x +y +6 =0

OpenStudy (anonymous):

y-10=2x-16 y-2x=-16+10 y-2x+6=0 -2x+y+6=0

OpenStudy (anonymous):

thanku

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

write y-6=-3(x+2)